OpenStudy (anonymous):

Given that d/dx 1-2x / square root (2+x^2) = g(x)/ square root (2+x^2)^3 ,find g(x).

terenzreignz (terenzreignz):

Let's just clear a few ambiguities... <ahem> Given that \[\Large \frac{d}{dx}\left[\frac{1-2x}{\sqrt{2+x^2}}\right]=\frac{g(x)}{(2+x^2)^3}\] find \(\large g(x)\) Is that it? :)

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

but must square root at the deneminator for RHS

terenzreignz (terenzreignz):

Well, the typical way I'd go about this is drop everything and just find the derivative of \[\Large \frac{1-2x}{\sqrt{2+x^2}}\] outright :D At least, first...

OpenStudy (anonymous):

my final answer for numerator is x + 4 i'm not sure whether there is a x

terenzreignz (terenzreignz):

Wait what? More like this? \[\Large \frac{g(x)}{\left[\sqrt{2+x^3}\right]^3}\]

OpenStudy (anonymous):

The cube belongs to the (2 + x^2) ^3/2

terenzreignz (terenzreignz):

Okay, much more acceptable :3 \[\Large \frac{g(x)}{(2+x^2)^{\frac32}}\] This, then?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

okay, well, let's go about differentiating \[\Large \frac{1-2x}{\sqrt{2+x^2}}\] Or have you already done that?

OpenStudy (anonymous):

i did that and my numerator i got x +4 is that right?

OpenStudy (anonymous):

I took out the negative sign in my numerator.

terenzreignz (terenzreignz):

Hang on...

terenzreignz (terenzreignz):

Oh. Well, you're asked what g(x) is, meaning the entire numerator, so you have to include the negative ... so it should be -x - 4